# On the multiplicity of weak solutions for a class of coupled quasilinear elliptic systems

### v3

@@ 4. The Concrete Palais-Smale condition
We recall an important compactness result, which follows by [20, Theorem 3.10].
@@
As consequence, we can state the following:
@@
For every $c\in\mathbb{R}$ the following facts are equivalent:
@@
$\int_{\Omega_{1}^{\rho}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}\leq\varepsilon\int_{\Omega_{2}^{\rho}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}+K(\rho,\varepsilon),$
$\Omega_{1}^{\rho}=\{x\in\Omega\ :\ |u_{n,k}|\leq\rho\},$ $\Omega_{2}^{\rho}=\{x\in\Omega\ :\ |u_{n,k}|>\rho\}.$
@@
We have that $\omega_{n,k}\to 0$ in $W^{-1,\boldsymbol{p}^{\prime}}(\Omega)$ as $n\to+\infty$ and for every $k=1,\dots,d$. We use an iterative procedure.
@@
Since $|\theta_{1,k}|\leq\sigma$ and $\{u_{n,j}\}$ is uniformly bounded in $L^{q_{j}\frac{q_{k}-1}{q_{k}}}(\Omega)$, we get
$\int_{\Omega}g_{k}(x,u_{n})\theta_{1,k}\leq\tilde{C},\qquad\tilde{C}=\tilde{C}(\sigma).$
@@
Notice that the left hand side is equal to
@@
We call $\theta(s)=(\theta_{1}(s),\dots,\theta_{d}(s))$. Hence,
@@
The first term in the right hand sign is uniformly bounded by a constant $c_{1}(l,\delta)$. We have also the following estimate by ($a.2$):
@@
Now, we consider the left hand side in (4.6). Hypothesis ($a.3$) implies that
@@
$\int_{\Omega_{1}^{\rho}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}$ $\leq\int_{I_{l\sigma}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}\leq K_{1}^{(l)}\int_{A_{l\sigma}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}+K_{2}^{(l)}$
@@
$\leq\frac{\nu\delta+\delta}{\nu-\delta}K_{1}^{(l)}\int_{\Omega_{2}^{\rho}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}+K_{3}(\rho,\delta),$
@@
$\int_{\Omega}D_{s}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}u_{n,k}$ $=\int_{\Omega_{1}^{R}}D_{s}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}u_{n,k}$
$\quad+\int_{\Omega_{2}^{R}}D_{s}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}u_{n,k}$
$\leq CR\int_{\Omega_{1}^{R}}|\nabla u_{n,k}|^{p_{k}}+\gamma\int_{\Omega_{2}^{R}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}},$
where the set $\Omega_{1}^{R},\Omega_{2}^{R}$ are defined in Proposition 4.3.
Let $\gamma^{\prime}\in(\gamma,\mu-\overline{p})$ and $\varepsilon>0$ such that $\frac{CR\varepsilon}{\nu}\leq\gamma^{\prime}-\gamma$. According to Proposition 4.3:
$\int_{\Omega}D_{s}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}u_{n,k}$ $\leq\frac{CR}{\nu}\int_{\Omega_{1}^{R}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}+\gamma\int_{\Omega_{2}^{R}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}$
$\leq\left(\frac{CR\varepsilon}{\nu}+\gamma\right)\int_{\Omega_{2}^{R}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}+K(R,\varepsilon)$
@@
$c\left(1+\|u_{n}\|_{W_{0}^{1,\boldsymbol{p}}}\right)$ $\geq f(u_{n})-\frac{1}{\mu}\langle\omega_{n},u_{n}\rangle=f(u_{n})-\frac{1}{\mu}\sum_{k=1}^{d}\langle\omega_{n,k},u_{n,k}\rangle$
@@ 5. Main Theorem
and
$|G(x,s)|\leq\kappa\sum_{k=1}^{d}\left(1+|s_{k}|^{q_{k}}\right).$
@@
We consider $\varphi_{h}=(\varphi_{h,p_{1}},\dots,\varphi_{h,p_{d}})\in W_{0}^{1,\boldsymbol{p}}(\Omega)\cap L^{\infty}(\Omega;\mathbb{R}^{d})$ where $\varphi_{h,p_{k}}$ is the eigenfunction of the problem:
$\begin{cases}-\Delta_{p_{k}}\varphi_{h,p_{k}}=\lambda_{h,p_{k}}\varphi_{h,p_{k}}&\text{in $\Omega$}\\ \varphi_{h,p_{k}}=0&\text{on $\partial\Omega$},\end{cases}$
and we define ${\mathcal{W}}_{h_{0}}:=\overline{\text{span}\{\varphi_{h}\ :h\geq h_{0}\}}$. We recall that $\lambda_{h,p_{k}}\to+\infty$ for every $k=1,\dots,d$ as $h\to+\infty$, see for instance [6, 18]. Furthermore, by [6, Proposition 5.4], we have that
(5.1) $\lambda_{h_{0},p_{k}}\int_{\Omega}|u_{k}|^{p_{k}}\leq\int_{\Omega}|\nabla u_{k}|^{p},\qquad\text{for every $u=(u_{1},\dots,u_{d})\in\mathcal{W}_{h_{0}}$.}$
Hence, for every $u\in\mathcal{W}_{h_{0}}$, applying (5.1) and Sobolev’s embedding yields
@@
Hence, setting $a_{k}=\|u_{k}\|_{W_{0}^{1,p_{k}}}$, we conclude that
@@
$\geq\frac{\nu}{\overline{p}}c_{1}\left(\sum_{k=1}^{d}\|u_{k}\|_{W_{0}^{1,p_{k}}}\right)^{\tilde{p}}-\left(\kappa|\Omega|+\frac{\nu}{\overline{p}}c_{2}\right)+o_{h_{0}}(1),$
where $o_{h_{0}}(1)\to 0\ \text{as $h_{0}\to+\infty$.}$ Then, $f(u)\geq\alpha>0$ on $\mathcal{W}_{h_{0}}\cap\partial B_{\rho}$ for a suitable choice of $\rho>0$ and for $h_{0}\geq 1$ large enough. Indeed, for every $u\in\partial B_{\rho}$, we have that
$f(u)\geq\frac{\nu}{\overline{p}}c_{1}\rho^{\tilde{p}}-\tilde{\kappa}+o_{h_{0}}(1)\geq 1+o_{h_{0}}(1),$
@@
Let $\varphi\in V\cap L^{\infty}(\Omega;\mathbb{R}^{d})$ with $\|\varphi\|=1$. By ($g.2$) we have that
$G(x,s)\geq b_{0}(x)|s|^{\mu}-a_{0}(x),\ \ \text{with $b_{0}\in L^{1}(\Omega),\ a_{0}\in L^{1}(\Omega)$},$
@@
(5.2) $f(t\varphi)\leq\sum_{k=1}^{d}\frac{Ct^{p_{k}}}{p_{k}}-t^{\mu}\int_{\Omega}b_{0}(x)|\varphi|^{\mu}-\|a_{0}\|_{L^{1}}\to-\infty,$
as $t\to+\infty$.
Since $W_{0}^{1,\boldsymbol{p}}(\Omega)\cap L^{\infty}(\Omega;\mathbb{R}^{d})$ is dense in $W_{0}^{1,\boldsymbol{p}}(\Omega)$, we have that for each $v\in V$ there exists $\{v_{n}\}\subset V\cap L^{\infty}(\Omega;\mathbb{R}^{d})$ with $v_{n}\to v$ in $W_{0}^{1,\boldsymbol{p}}$. Therefore,
$f(tv)=\lim_{n\to\infty}f(tv_{n})\leq 0,$
for $t>0$ large enough by (5.2).
@@
Let $c\in\mathbb{R}$ and let $\{u_{n}\}\subset W_{0}^{1,\boldsymbol{p}}(\Omega)$ be a $(PS)$-sequence at level $c$. According to Corollary 2.8-$(ii)$, $\{u_{n}\}$ is also a $(CPS)$-sequence and Theorems 4.4 and 4.2 imply that the $(CPS)$-condition holds. Corollary 2.8-$(iii)$ allow us to conclude that the functional satisfies also the $(PS)$-condition.
Now, according to Theorem 2.5 we find a sequence of critical points $\{u_{n}\}\subset W_{0}^{1,\boldsymbol{p}}(\Omega)$ for $f$ such that $f(u_{n})\to+\infty$ and Theorem 3.1 proves that $u_{n}\in W_{0}^{1,\boldsymbol{p}}(\Omega)\cap L^{\infty}(\Omega;\mathbb{R}^{d})$, concluding the proof. ∎

### v4

@@ 4. The Concrete Palais-Smale condition
We recall a significant compactness result, which follows from [20, Theorem 3.10].
@@
As a consequence, we can state the following:
@@
For every $c\in\mathbb{R}$, the following facts are equivalent:
@@
$\int_{\Omega_{n,\rho}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}\leq\varepsilon\int_{\Omega_{n,\rho}^{c}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}+K(\rho,\varepsilon),$
$\Omega_{n,\rho}=\{x\in\Omega\ :\ |u_{n,k}|\leq\rho\},$ $\Omega_{n,\rho}^{c}=\{x\in\Omega\ :\ |u_{n,k}|>\rho\}.$
@@
We have that $\omega_{n,k}\to 0$ in $W^{-1,p_{k}^{\prime}}(\Omega)$ as $n\to+\infty$ and for every $k=1,\dots,d$. We use an iterative procedure.
@@
Since $|\theta_{1,k}|\leq\sigma$ and $\{u_{n,j}\}$ is uniformly bounded in $L^{q_{j}\frac{q_{k}-1}{q_{k}}}(\Omega)$ for every $j=1,\dots,d$, we get
$\int_{\Omega}g_{k}(x,u_{n})\theta_{1,k}\leq\tilde{C},\qquad\text{for some }\tilde{C}=\tilde{C}(\sigma)>0.$
@@
Notice that the left-hand side is equal to
@@
We denote $\theta(s)=(\theta_{1}(s),\dots,\theta_{d}(s))$. Hence,
@@
The first term in the right-hand sign is uniformly bounded by a positive constant $c_{1}(l,\delta)$. We also have the following estimate by ($a.2$):
@@
Now, we consider the left-hand side in (4.6). Hypothesis ($a.3$) implies that
@@
$\int_{\Omega_{n,\rho}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}$ $\leq\int_{I_{l\sigma}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}\leq K_{1}^{(l)}\int_{A_{l\sigma}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}+K_{2}^{(l)}$
@@
$\leq\frac{\nu\delta+\delta}{\nu-\delta}K_{1}^{(l)}\int_{\Omega_{n,\rho}^{c}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}+K_{3}(\rho,\delta),$
@@
$\int_{\Omega}D_{s}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}u_{n,k}$ $=\int_{\Omega_{n,R}}D_{s}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}u_{n,k}$
$\quad+\int_{\Omega_{n,R}^{c}}D_{s}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}u_{n,k}$
$\leq C_{0}R\int_{\Omega_{n,R}}|\nabla u_{n,k}|^{p_{k}}+\gamma\int_{\Omega_{n,R}^{c}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}},$
where the sets $\Omega_{n,R},\Omega_{n,R}^{c}$ are defined in Proposition 4.3.
Let $\gamma^{\prime}\in(\gamma,\mu-\overline{p})$ and $\varepsilon>0$ such that $\frac{C_{0}R\varepsilon}{\nu}\leq\gamma^{\prime}-\gamma$. According to Proposition 4.3:
$\int_{\Omega}D_{s}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}u_{n,k}$ $\leq\frac{C_{0}R}{\nu}\int_{\Omega_{n,R}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}+\gamma\int_{\Omega_{n,R}^{c}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}$
$\leq\left(\frac{C_{0}R\varepsilon}{\nu}+\gamma\right)\int_{\Omega_{n,R}^{c}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}+K(R,\varepsilon)$
@@
$C+1+\|u_{n}\|_{W_{0}^{1,\boldsymbol{p}}}$ $\geq f(u_{n})-\frac{1}{\mu}\langle\omega_{n},u_{n}\rangle=f(u_{n})-\frac{1}{\mu}\sum_{k=1}^{d}\langle\omega_{n,k},u_{n,k}\rangle$
@@ 5. Main Theorem
We set $\sigma_{j,k}=q_{j}\frac{q_{k}-1}{q_{k}}$. Exploiting Young’s inequality, we obtain that
$|s_{j}|^{q_{j}\frac{q_{k}-1}{q_{k}}}|s_{k}|\leq\kappa_{1}|s_{j}|^{q_{j}}+\kappa_{2}|s_{k}|^{q_{k}},\qquad|s_{k}|\leq\kappa_{3}+\kappa_{4}|s_{k}|^{q_{k}},$
with $\kappa_{1},\kappa_{2},\kappa_{3},\kappa_{4}>0$, and
$|G(x,s)|\leq\kappa\sum_{k=1}^{d}\left(1+|s_{k}|^{q_{k}}\right),$
for some $\kappa>0$.
@@
According to [7, Section 5], there exist two sequences $\{\lambda_{h,p_{k}}\}_{h\in\mathbb{N}}\subset(0,+\infty)$ and $\{\varphi_{h,p_{k}}\}_{h\in\mathbb{N}}\subset W_{0}^{1,p_{k}}(\Omega)$ such that
$\begin{cases}-\Delta_{p_{k}}\varphi_{h,p_{k}}=\lambda_{h,p_{k}}|\varphi_{h,p_{k}}|^{p_{k}-2}\varphi_{h,p_{k}}&\text{in $\Omega$},\\ \varphi_{h,p_{k}}=0&\text{on $\partial\Omega$},\end{cases}$
and
$W_{0}^{1,p_{k}}(\Omega)=\operatorname{span}\{\varphi_{1,p_{k}},\dots,\varphi_{h_{0}-1,p_{k}}\}\oplus\mathcal{W}_{h_{0},p_{k}},$
where $\mathcal{W}_{h_{0},p_{k}}\subset W_{0}^{1,p_{k}}(\Omega)$ is a subspace of finite codimension.
Furthermore, $\lambda_{h,p_{k}}\to+\infty$ for every $k=1,\dots,d$ as $h\to+\infty$, and
(5.1) $\lambda_{h_{0},p_{k}}\int_{\Omega}|u_{k}|^{p_{k}}\leq\int_{\Omega}|\nabla u_{k}|^{p_{k}},\qquad\text{for every }u_{k}\in\mathcal{W}_{h_{0},p_{k}},\ k=1,\dots,d.$
Finally, we define
$\mathcal{W}_{h_{0}}:=\mathcal{W}_{h_{0},p_{1}}\times\dots\times\mathcal{W}_{h_{0},p_{d}}.$
Hence, for every $u\in\mathcal{W}_{h_{0}}$, (5.1) and Sobolev’s embedding yield
@@
Hence, setting $a_{k}=\|u_{k}\|_{W_{0}^{1,p_{k}}}$, and recalling that $\overline{p}=\max\{p_{1},\dots,p_{d}\}$, we conclude that
@@
$\geq\frac{\nu}{\overline{p}}c_{1}\left(\sum_{k=1}^{d}\|u_{k}\|_{W_{0}^{1,p_{k}}}\right)^{\tilde{p}}-\left(\kappa|\Omega|+\frac{\nu}{\overline{p}}c_{2}\right)-\sum_{k=1}^{d}\lambda_{h_{0},p_{k}}^{-\frac{\tau_{k}}{p_{k}}}\|u_{k}\|_{W_{0}^{1,p_{k}}}^{q_{k}},$
where
$\sum_{k=1}^{d}\lambda_{h_{0},p_{k}}^{-\frac{\tau_{k}}{p_{k}}}\|u_{k}\|_{W_{0}^{1,p_{k}}}^{q_{k}}\to 0\qquad\text{as $h_{0}\to+\infty$.}$
Then, $f(u)\geq\alpha>0$ on $\mathcal{W}_{h_{0}}\cap\partial B_{\rho}$ for a suitable choice of $\rho,\alpha>0$ and for $h_{0}\geq 1$ large enough. Indeed, for every $u\in\partial B_{\rho}$, we have that there exists $\tilde{\kappa}>0$ such that
$f(u)\geq\frac{\nu}{\overline{p}}c_{1}\rho^{\tilde{p}}-\tilde{\kappa}-\sum_{k=1}^{d}\lambda_{h_{0},p_{k}}^{-\frac{\tau_{k}}{p_{k}}}\|u_{k}\|_{W_{0}^{1,p_{k}}}^{q_{k}}\geq 1-\sum_{k=1}^{d}\lambda_{h_{0},p_{k}}^{-\frac{\tau_{k}}{p_{k}}}\|u_{k}\|_{W_{0}^{1,p_{k}}}^{q_{k}},$
@@
Let $\varphi\in V$ with $\|\varphi\|=1$. By ($g.2$) we have that
$G(x,s)\geq b_{0}(x)|s|^{\mu}-a_{0}(x),\ \ \text{with $b_{0}\in L^{\infty}(\Omega),\ a_{0}\in L^{1}(\Omega)$},$
@@
(5.2) $f(t\varphi)\leq\sum_{k=1}^{d}\frac{Ct^{p_{k}}}{p_{k}}-t^{\mu}\int_{\Omega}b_{0}(x)|\varphi|^{\mu}+\|a_{0}\|_{L^{1}}\to-\infty,$
as $t\to+\infty$. Therefore, the thesis follows by taking $R>0$ large enough.
@@
Let $c\in\mathbb{R}$ and let $\{u_{n}\}\subset W_{0}^{1,\boldsymbol{p}}(\Omega)$ be a $(PS)$-sequence at level $c$. According to Corollary 2.8-$(ii)$, $\{u_{n}\}$ is also a $(CPS)$-sequence and Theorems 4.4 and 4.2 imply that the $(CPS)$-condition holds. Corollary 2.8-$(iii)$ allows us to conclude that the functional also satisfies the $(PS)$-condition.
Now, according to Theorem 2.5 we find a sequence of critical points $\{u_{n}\}\subset W_{0}^{1,\boldsymbol{p}}(\Omega)$ for $f$ such that $f(u_{n})\to+\infty$, as $n\to+\infty$, and Theorem 3.1 proves that $u_{n}\in W_{0}^{1,\boldsymbol{p}}(\Omega)\cap L^{\infty}(\Omega;\mathbb{R}^{d})$, concluding the proof. ∎